# QCD axion dark matter in the dark dimension

### v1

@@ 3.2 Enhanced axion abundance through resonant conversion
In this subsection, we investigate the enhancement of the QCD axion DM abundance within the context of a simple two-axion mixing scenario. In particular, we consider the mixing between one QCD axion and one ALP.
The low-energy effective Lagrangian that describes this two-axion mixing can be formulated as follows
@@
where $\Theta=\varphi/f_{A}$ is the ALP angle, $\varphi$ represents the ALP field, $m_{A}$ and $f_{A}$ denote the ALP mass and decay constant, respectively, and $n_{ij}$ are the domain wall numbers. Notice that the ALP field is considered as the simplest case with a constant mass $m_{A}$. For our purpose, here the domain wall numbers should be taken as [^1] The mixing potential for this particular selection was first explored in ref. Cyncynates:2023esj , where it was used to study the mixing between the QCD axion and sterile axion, ultimately resulting in the heavy QCD axion as DM. Additionally, for a different selections of domain wall numbers, the axion mixing potential can lead to a suppression of the QCD axion abundance, see $\rm e.g.$ refs. Kitajima:2014xla ; Daido:2015cba ; Ho:2018qur ; Li:2023xkn ; Li:2023uvt ; Li:2024okl ; Murai:2024nsp .
@@
which represent the two distinct axion states arising from the mixing process. These eigenstates are associated with the mass eigenvalues $m_{h}(T)$ and $m_{l}(T)$, respectively.
Figure 1: The temperature $T_{R}$ as a function of the axion mass ratio $m_{a,0}/m_{A}$. Here we set $f_{a}=1.0\times 10^{10}\,\rm GeV$. The red, blue, and purple solid lines represent scenarios with $f_{A}=1.1\times 10^{10}\,\rm GeV$, $2.0\times 10^{10}\,\rm GeV$, and $4.0\times 10^{10}\,\rm GeV$, respectively. The black dashed line corresponds to the case where $f_{A}$ is much greater than $f_{a}$. The gray dashed line represents the QCD phase transition critical temperature $T_{\rm QCD}$.
Next we discuss the resonant conversion within the context of axion mixing. For our analysis, we consider the general prerequisites for axion resonant conversion, which are expressed as follows
@@
See figure 1 for the distribution of $T_{R}$ as a function of the axion mass ratio $m_{a,0}/m_{A}$ with the QCD axion decay constant $f_{a}=1.0\times 10^{10}\,\rm GeV$. This figure highlights the dramatic changes in $T_{R}$ as the ALP decay constant $f_{A}$ increases. We can observe that $T_{R}$ demonstrates an asymptotic behavior when $f_{A}$ is slightly greater than $f_{a}$, and it is not necessary for $f_{A}$ to be significantly larger than $f_{a}$ for this distribution to manifest.
At temperatures above $T_{R}$, the heavy mass eigenstate $a_{h}$ primarily consists of the ALP, while the light mass eigenstate $a_{l}$ is dominated by the QCD axion. As the cosmic temperature decreases, the mass eigenvalues $m_{h}(T)$ and $m_{l}(T)$ converge at $T=T_{R}$ and then diverge. Below the temperature $T_{R}$, $a_{h}$ becomes the QCD axion, and $a_{l}$ transforms into the ALP. Here, the axion energy transition at $T_{R}$ is presumed to be adiabatic, which is approximately valid when
$T_{i,a}\gg T_{R}\,,$ (57)
where $T_{i,a}$ represents the QCD axion oscillation temperature. In this case, the ALP begins oscillating prior to $T_{i,a}$. Subsequently, it is crucial to examine the range of $m_{A}$, which can be characterized by
@@
$\rho_{A,i}=\frac{1}{2}m_{A}^{2}f_{A}^{2}\Theta_{i}^{2}\,,$ (60)
where $T_{i,A}$ is given by $3H(T_{i,A})=m_{A}$. For temperatures $T_{R}<T<T_{i,A}$, the ALP energy density is adiabatic invariant with the comoving number $N_{A}\equiv\rho_{A}a^{3}/m_{A}$ and it at $T_{R}$ is expressed as
$\rho_{A,R}=\frac{1}{2}m_{A}^{2}f_{A}^{2}\Theta_{i}^{2}\left(\frac{a_{i,A}}{a_{R}}\right)^{3}\,,$ (61)
where $a_{i,A}$ and $a_{R}$ are the scale factors at $T_{i,A}$ and $T_{R}$, respectively. At $T=T_{R}$, the energy density of the ALP, $\rho_{A,R}$, is transferred to the QCD axion, $\rho_{a,R}$. For temperatures below $T_{R}$, the QCD axion energy density is adiabatic invariant and it at $T_{0}$ is given by
$\rho^{\prime}_{a,0}=\frac{1}{2}m_{a,0}m_{A}f_{A}^{2}\Theta_{i}^{2}\left(\frac{a_{i,A}}{a_{0}}\right)^{3}\,,$ (62)
where $T_{0}$ is the present CMB temperature, and $a_{0}$ is the scale factor at $T_{0}$. Comparing this with the case without mixing, we find that the QCD axion energy density can be modified by a factor
@@
Figure 2: The ALP constraints in the $\{m_{A},1/f_{A}\}$ plane. The red and blue solid lines correspond to $f_{a}=10^{9}\,\rm GeV$ and $10^{10}\,\rm GeV$, respectively. The shadow area between these two lines represents the allowed ALP parameter space to account for the overall DM abundance in our scenario. Additional constraints on ALP within this plane are sourced from ref. ciaran_o_hare_2020_3932430 .
@@
$\Omega^{\prime}_{a}h^{2}\simeq 0.17\left(\dfrac{f_{a}}{10^{10}\,\rm GeV}\right)^{-1}\left(\dfrac{f_{A}}{10^{12}\,\rm GeV}\right)^{2}\left(\dfrac{m_{A}}{10^{-5}\,\rm eV}\right)^{-1/2}\left\langle\Theta_{i}^{2}f(\Theta_{i})\right\rangle\,,$ (65)
where we neglect some trivial factors. By assuming a typical ALP mass of $m_{A}=10^{-5}\,\rm eV$, and considering the observed cold DM abundance $\Omega_{\rm DM}h^{2}\simeq 0.12$ with the QCD axion in the dark dimension, where $f_{a}\sim 10^{9}-10^{10}\,\rm GeV$, we find that the ALP decay constant falls within the range
@@
Utilizing eqs. (59) and (65), we also show in figure 2 the allowed ALP parameter space in the $\{m_{A},1/f_{A}\}$ plane. The red and blue solid lines depict the scenarios where $f_{a}=10^{9}\,\rm GeV$ and $10^{10}\,\rm GeV$, respectively. The shaded gray quadrilateral area bounded by these two lines represents the allowed ALP parameter space within our scenario, encompassing the range that accounts for the overall DM abundance via resonant conversion. We also find that a significant portion of ALPs falling within this specified range have the potential to be detected by future experiments.
Additionally, in the context of considering a scenario with multiple QCD axions mass mixing Li:2024okl , the abundance of QCD axion DM can also be effectively enhanced, depending on the relationship among the QCD axion decay constants. That scenario should also be capable of explaining the issue of QCD axion DM abundance in the dark dimension, requiring further investigation.
@@ 4 Conclusion
In summary, we have investigated the issue of QCD axion DM within the context of the recently proposed dark dimension scenario, and demonstrated that the QCD axion can account for the overall DM abundance through a simple axion mixing mechanism. We first provide a concise overview of the QCD axion in the dark dimension scenario. Subsequently, we discuss the QCD axion DM in this scenario through the misalignment mechanism. Then we discuss the enhancement of the QCD axion DM abundance within the context of a simple two-axion mixing scenario.
The dark dimension scenario is motivated by the minuscule nature of dark energy, informed by Swampland principles and constrained by observational data. In this context, the axions can be localized on the SM brane. By applying the WGC to the QCD axion, one can derive an upper bound for the axion decay constant, $f_{a}\lesssim M_{5}\sim 10^{9}-10^{10}\,\rm GeV$, where $M_{5}$ is the five-dimensional Planck mass. When combined with observational lower bounds, this suggests that $f_{a}$ falls within the range $f_{a}\sim 10^{9}-10^{10}\,\rm GeV$, corresponding to an axion mass $m_{a}\sim 10^{-3}-10^{-2}\,\rm eV$. According to the misalignment mechanism, the QCD axion at this scale could constitute a minor fraction of the total DM density $\sim 10^{-3}-10^{-2}$ if without the fine-tuning of the initial misalignment angle. In order to enhance the QCD axion DM abundance, we consider a straightforward two-axion mixing mechanism, specifically the resonant conversion of an ALP into a QCD axion. Our findings suggest that, in a scenario where the ALP possesses a mass of approximately $m_{A}\sim 10^{-5}\,\rm eV$ and a decay constant of $f_{A}\sim 10^{11}\,\rm GeV$, the QCD axion in the dark dimension scenario can account for the overall DM abundance. This is supported by a detailed analysis of the axion mixing potential and the temperature effects on the axion energy density. Furthermore, future experiments aimed at detecting ALPs within this specific range have the potential to further validate this scenario. Additionally, a scenario with multiple QCD axions mass mixing should also be able to explain the issue discussed in this work. Our work highlights the potential of the QCD axion within the dark dimension scenario to account for the DM.
@@ Acknowledgments
We thank Yu-Feng Zhou for useful discussions. This work was supported by the Key Laboratory of Theoretical Physics in Institute of Theoretical Physics, CAS.

### v2

@@ 3.2 Enhanced axion abundance through resonant conversion
In this subsection, we investigate the enhancement of the QCD axion DM abundance within the context of a simple two-axion mixing scenario. In particular, we consider the mixing between one QCD axion and one ALP. The low-energy effective Lagrangian that describes this two-axion mixing can be formulated as follows
@@
where $\Theta=\varphi/f_{A}$ is the ALP angle, $\varphi$ represents the ALP field, $m_{A}$ and $f_{A}$ denote the ALP mass and decay constant, respectively, and $n_{ij}$ are the domain wall numbers. Notice that the single-field ALP is considered as the simplest case with a constant mass $m_{A}$. For our purpose, here the domain wall numbers should be taken as [^1] The mixing potential for this particular selection was first explored in ref. Cyncynates:2023esj , where it was used to study the mixing between the QCD axion and sterile axion, ultimately resulting in the heavy QCD axion as DM. Additionally, for different selections of domain wall numbers, the axion mixing potential can lead to a suppression of the QCD axion abundance, see $\rm e.g.$ refs. Kitajima:2014xla ; Daido:2015cba ; Ho:2018qur ; Li:2023det ; Li:2023xkn ; Li:2023uvt ; Murai:2024nsp .
@@
which represent the two distinct axion states arising from the mixing process. These eigenstates are associated with the mass eigenvalues $m_{h,l}(T)$, respectively.
Next we discuss the axion resonant conversion within the context of this specific mixing. For our analysis, we consider the general prerequisites for resonant conversion, which are expressed as follows
@@
See figure 1 for the distribution of $T_{R}$ as a function of the axion mass ratio $m_{a,0}/m_{A}$, with the QCD axion decay constant set at $f_{a}=1.0\times 10^{10}\,\rm GeV$. This figure highlights the substantial variations in $T_{R}$ as the ALP decay constant $f_{A}$ increases. Notably, we can observe that the resonant conversion temperature demonstrates an asymptotic behavior when $f_{A}$ is marginally greater than $f_{a}$, indicating that a significant disparity between $f_{A}$ and $f_{a}$ is not prerequisite for this distribution to emerge.
Figure 1: The resonant conversion temperature $T_{R}$ as a function of the axion mass ratio $m_{a,0}/m_{A}$. Here we set $f_{a}=1.0\times 10^{10}\,\rm GeV$. The red, blue, and purple solid lines represent scenarios with $f_{A}=1.1\times 10^{10}\,\rm GeV$, $2.0\times 10^{10}\,\rm GeV$, and $4.0\times 10^{10}\,\rm GeV$, respectively. Notice that $f_{A}$ cannot be equal to $f_{a}$ here. The black dashed line corresponds to the case where $f_{A}$ is much greater than $f_{a}$. The gray dashed line represents the QCD phase transition critical temperature $T_{\rm QCD}$.
At temperatures above $T_{R}$, the heavy mass eigenstate $a_{h}$ primarily consists of the ALP, while the light mass eigenstate $a_{l}$ is dominated by the QCD axion. As the cosmic temperature decreases, the mass eigenvalues $m_{h}(T)$ and $m_{l}(T)$ converge at $T=T_{R}$ and then diverge. Below the temperature $T_{R}$, $a_{h}$ becomes the QCD axion, and $a_{l}$ transforms into the ALP. Here, the axion energy transition at $T_{R}$ is presumed to be adiabatic, which is approximately valid when Cyncynates:2023esj
$T_{R}\ll T_{i,a}\,,$ (57)
where $T_{i,a}$ represents the QCD axion oscillation temperature shown in eq. (32). In this case, the ALP begins oscillating prior to $T_{i,a}$. Subsequently, it is crucial to examine the range of the ALP mass $m_{A}$, which can be characterized by
@@
To satisfy both sub-formulas simultaneously, we will choose a typical value of $m_{A}=10^{-5}\,\rm eV$ for the ALP mass in our subsequent discussion. Since the ALP mass is relatively flexible, this choice is reasonable.
@@
$\rho_{A,i}=\frac{1}{2}m_{A}^{2}f_{A}^{2}\left\langle\Theta_{i}^{2}f(\Theta_{i})\right\rangle\,,$ (60)
where $T_{i,A}$ is given by $3H(T_{i,A})=m_{A}$. For temperatures $T_{R}<T<T_{i,A}$, the ALP energy density is adiabatic invariant with the comoving number $N_{A}\equiv\rho_{A}\mathfrak{a}^{3}/m_{A}$. At $T_{R}$, this energy density is expressed as
$\rho_{A,R}=\frac{1}{2}m_{A}^{2}f_{A}^{2}\left\langle\Theta_{i}^{2}f(\Theta_{i})\right\rangle\left(\frac{\mathfrak{a}_{i,A}}{\mathfrak{a}_{R}}\right)^{3}\,,$ (61)
where $\mathfrak{a}_{i,A}$ and $\mathfrak{a}_{R}$ are the scale factors at $T_{i,A}$ and $T_{R}$, respectively. Then the energy density of the ALP, $\rho_{A,R}$, is transferred to the QCD axion, $\rho_{a,R}$. For temperatures below $T_{R}$, the QCD axion energy density is adiabatic invariant. At $T_{0}$, this energy density is given by
$\rho^{\prime}_{a,0}=\frac{1}{2}m_{a,0}m_{A}f_{A}^{2}\left\langle\Theta_{i}^{2}f(\Theta_{i})\right\rangle\left(\frac{\mathfrak{a}_{i,A}}{\mathfrak{a}_{0}}\right)^{3}\,,$ (62)
where $T_{0}$ is the present CMB temperature, and $\mathfrak{a}_{0}$ is the scale factor at $T_{0}$. Comparing this with the case without mixing, we find that the QCD axion energy density can be modified by a factor
@@
[…]
By assuming a typical ALP mass of $m_{A}=10^{-5}\,\rm eV$, and considering the observed cold DM abundance $\Omega_{\rm DM}h^{2}\simeq 0.12$ with the QCD axion in the dark dimension, where $f_{a}\sim 10^{9}-10^{10}\,\rm GeV$, we find that the ALP decay constant falls within the range
@@
Utilizing eqs. (59) and (65), we also show in figure 2 the allowed ALP parameter space in the $\{m_{A},1/f_{A}\}$ plane. The red and blue solid lines represent the scenarios where $f_{a}=10^{9}\,\rm GeV$ and $10^{10}\,\rm GeV$, respectively. The shaded gray quadrilateral area bounded by these two lines represents the allowed ALP parameter space within our scenario, encompassing the range that accounts for the overall DM abundance via axion resonant conversion.
Several aspects require further clarification and elaboration for a clear understanding. Regarding eq. (59), there are two key points to clarify. Firstly, it is presented considering that in certain regions, the ALP will no longer be lighter than the QCD axion. Secondly, we have considered keeping the resonant conversion temperature as low as possible compared to the axion oscillation temperature, to ensure that adiabatic conversion can occur completely. [^2] We have assumed the approximation that all ALP density is transferred to the QCD axion, which is acceptable for large disparities in axion decay constants. When the adiabatic condition is violated, this resonant conversion does not occur completely, and thus the allowed parameter space for ALP may expand. Regarding the extent to which adiabatic condition is violated, this remains uncertain and requires further complex numerical analysis. At least, this condition should not be violated in this work, since $f_{a}/f_{A}\sim 10^{-3}-10^{-2}\ll 1$. Nevertheless, the constraints presented here conservatively determine the ALP parameter space in the case of complete conversion. These two points are addressed in eqs. (58) and (57), which provide the mass range of ALP in eq. (59). Regarding eq. (65), it involves three variables: $f_{a}$, $f_{A}$, and $m_{A}$. To illustrate the ALP parameter space in the $\{m_{A},1/f_{A}\}$ plane, we need to fix the value of $f_{a}$. Consequently, there will be two lines corresponding to $f_{a}=10^{9}\,\rm GeV$ and $10^{10}\,\rm GeV$, represented as the red and blue lines in the figure, respectively. Naturally, the range of these red and blue lines is also constrained by eq. (59). Then the shaded area between these two lines represents the allowed region for the ALP that we aim to obtain in this work. Furthermore, we also find that a significant portion of ALPs falling within this specific range have the potential to be detected by future axion experiments.
Figure 2: The ALP constraints in the $\{m_{A},1/f_{A}\}$ plane. The red and blue solid lines correspond to the scenarios where the QCD axion decay constant is $f_{a}=10^{9}\,\rm GeV$ and $10^{10}\,\rm GeV$, respectively. The shaded area between these two lines represents the allowed ALP parameter space that can account for the overall DM abundance in our scenario. Additional constraints on ALP within this plane are sourced from ref. ciaran_o_hare_2020_3932430 .
Additionally, in the context of considering a scenario with multiple QCD axions mass mixing Li:2024okl , the abundance of QCD axion DM can also be effectively enhanced, depending on the relationship among axion decay constants. This scenario may also provide an explanation for the issue of QCD axion DM abundance in the dark dimension scenario, requiring further investigation.
@@ 3.2.1 Brief discussion on adiabatic condition
3.3 Possible origins of ALP with a higher-scale decay constant
In this subsection, we discuss the possible origins of the ALP with a higher-scale decay constant that is required to satisfy the preceding eq. (66). We speculate that the ALP, which falls within this specific range, may originate from the grand unification of gauge forces in the context of the dark dimension, where realizing such unification is highly constrained Heckman:2024trz . In this scenario, the WGC constraint on the ALP decay constant is modified by a factor of $M_{5}^{1/2}L_{g}^{1/2}\sim 10^{3}$, resulting in
$f_{A}\lesssim 10^{3}\times M_{5}\sim 10^{12}-10^{13}\,\rm GeV\,,$ (67)
which sits comfortably in the permitted experimental range, where $L_{g}$ represents the diameter of the brane accommodating gauge fields,
$L_{g}\lesssim\left(1-10\,{\rm TeV}\right)^{-1}\,.$ (68)
However, in this case, the QCD axion may also be subject to this constraint, leading to a larger decay constant too. This, in turn, may reduce the need for the ALP resonant conversion discussed here, although having a more weakly-coupled axion might be less appealing. Unless, of course, the decay constant of the QCD axion in this scenario is also subject to other unknown constraints.
Specifically, it is natural for us to further consider the framework of what is termed the “dark-dimensional axiverse,"""" where both the QCD axion and a large number of ALPs coexist, leading to the Lagrangian
$\mathcal{L}\supset\dfrac{1}{2}\sum_{i}f_{i}^{2}\left(\partial a_{i}\right)^{2}-\sum_{i}\Lambda_{i}^{4}\left[1-\cos\left(\sum_{j}n_{ij}a_{j}\right)\right]\,,$ (69)
where $a_{i}$, $f_{i}$, $\Lambda_{i}$, and $n_{ij}$ represent the axion angles, decay constants, overall scales, and domain wall numbers, respectively. Within such a model, the decay constant of the QCD axion can be set at a relatively small value, whereas the decay constant of some ALPs may be significantly enhanced. Notice also that this type of multi-axion scenario, with hierarchical axion decay constants, can be realized in string theory, such as the commonly encountered type IIB string axiverse Cicoli:2012sz . Nevertheless, this exceeds the scope of the present paper and merits further exploration in future work.
@@ 4 Conclusion
In summary, we have investigated the issue of QCD axion DM within the context of the recently proposed dark dimension scenario, and demonstrated that the QCD axion can account for the overall DM abundance through the axion mass mixing mechanism. We first provide a concise overview of the QCD axion in the dark dimension. Subsequently, we discuss the QCD axion DM in this scenario through the misalignment mechanism. Then we discuss the enhancement of the axion DM abundance within the context of a simple two-axion mixing scenario.
The dark dimension scenario is motivated by the minuscule nature of dark energy, informed by Swampland principles and constrained by observational data. In this context, axions can be localized on the SM brane. By applying the WGC to the QCD axion, one can derive an upper bound for the axion decay constant, $f_{a}\lesssim M_{5}\sim 10^{9}-10^{10}\,\rm GeV$, where $M_{5}$ is the five-dimensional Planck mass. When combined with observational lower bounds, this suggests that $f_{a}$ falls within the range $f_{a}\sim 10^{9}-10^{10}\,\rm GeV$, corresponding to an axion mass $m_{a}\sim 10^{-3}-10^{-2}\,\rm eV$. According to the misalignment mechanism, the QCD axion at this scale could constitute a minor fraction of the total DM density $\sim 10^{-3}-10^{-2}$ if without the fine-tuning of the initial misalignment angle. In order to enhance the QCD axion DM abundance, we consider a straightforward two-axion mixing mechanism, specifically the resonant conversion of an ALP into the QCD axion. Our findings suggest that, in a scenario where the ALP possesses a mass of approximately $m_{A}\sim 10^{-5}\,\rm eV$ and a decay constant of $f_{A}\sim 10^{11}\,\rm GeV$, the QCD axion in the dark dimension scenario can account for the overall DM abundance. This is supported by a detailed analysis of the axion mixing potential and the temperature effects on the axion energy density. We also speculate that the ALP required within this specific range of decay constant may stem from the grand unification in the context of the dark dimension. Furthermore, future experiments aimed at detecting ALPs within this range have the potential to further validate this scenario. Additionally, a scenario with multiple QCD axions mass mixing should also be able to explain the issue discussed in this work. Our work highlights the potential of the QCD axion within the dark dimension scenario to account for the cold DM abundance.
@@ Acknowledgments
We thank Michele Cicoli and Yu-Feng Zhou for useful discussions. This work was supported by the Key Laboratory of Theoretical Physics in Institute of Theoretical Physics, CAS.